OpenStudy (anonymous):

find the exact value?! sin [arc cos(15/17)]

OpenStudy (anonymous):

Use a triangle to understand this:  Let \[\huge \theta=\cos^{-1}\frac{15}{17}\]

OpenStudy (anonymous):

Find the question mark and then you can find what sintheta is using your trig knowledge.

OpenStudy (anonymous):

well the missing side is 8 where do i go from there??

OpenStudy (anonymous):

Then, you know that sintheta=opposite/hypotenuse

OpenStudy (anonymous):

So what would sintheta=? When you know that the opposite side is 8.